ariannaayala21:

How much compound interest will $50,000 earn in 10 years at 3.4% APR compounded yearly, providing no deposits or withdrawals are made? Hint: Use the formula A = P(1.00 + r)n.

Vocaloid:

P = 50,000 r = 0.034 (3.4% as a decimal) n = 10 years

Vocaloid:

A = P(1.00 + r^)n = ?